Here is another one I can not figure out. Please help.
In exercise 36, if f(x)=5x-1, find the following:
f(a-2)

Answer by ankor@dixie-net.com(12687)

if f(x)=5x-1, find the following:
f(a-2)
:
Just replace x with (a-2):
f(a-2) = 5(a-2) - 1
f(a-2) = 5a - 10 -1
f(a-2) = 5a - 11